OpenStudy (anonymous):

is (-2,8) a solution of 3x-4y=18??

OpenStudy (amistre64):

plug it in to test the case

OpenStudy (anonymous):

Plug in your point for x and y and see if it's equals

OpenStudy (amistre64):

(x=-2,y=8) 3x-4y=18 -2 8 ----------- -6-32=18

OpenStudy (anonymous):

substitute: 3(-2) - 4(8) = -6 - 32 = -38 which does not equal 18

OpenStudy (anonymous):

thanks